### v1

# TOPA: Extend Large Language Models for Video Understanding via Text-Only Pre-Alignment

Abstract
Recent advancements in image understanding have benefited from the extensive use of web image-text pairs. However, video understanding remains a challenge despite the availability of substantial web video-text data. This difficulty primarily arises from the inherent complexity of videos and the inefficient language supervision in recent web-collected video-text datasets. In this paper, we introduce Text-Only Pre-Alignment (TOPA), a novel approach to extend large language models (LLMs) for video understanding, without the need for pre-training on real video data. Specifically, we first employ an advanced LLM to automatically generate Textual Videos comprising continuous textual frames, along with corresponding annotations to simulate real video-text data. Then, these annotated textual videos are used to pre-align a language-only LLM with the video modality. To bridge the gap between textual and real videos, we employ the CLIP model as the feature extractor to align image and text modalities. During text-only pre-alignment, the continuous textual frames, encoded as a sequence of CLIP text features, are analogous to continuous CLIP image features, thus aligning the LLM with real video representation. Extensive experiments, including zero-shot evaluation and finetuning on various video understanding tasks, demonstrate that TOPA is an effective and efficient framework for aligning video content with LLMs. In particular, without training on any video data, the TOPA-Llama2-13B model achieves a Top-1 accuracy of 51.0% on the challenging long-form video understanding benchmark, Egoschema. This performance surpasses previous video-text pre-training approaches and proves competitive with recent GPT-3.5-based video agents.
1 Introduction
Image-language understanding has made large advancements in both image-language alignment [22, 43] and Multimodal Large Language Models (MLLMs) [1, 23, 29, 79], benefiting from pre-training on large-scale noise-paired image-text data collected from the web [6, 17, 45, 47, 46]. This raises a question: Can we mirror this success in video-language understanding? Research [40, 56, 67, 76] has explored pretraining video-language models on millions of web video-text data [3, 35], achieving promising results in basic video tasks such as video-text retrieval, video captioning, and video question answering across conventional video benchmarks. However, recent research reveals that these models struggle with the challenging long-form video understanding benchmark, i.e., Egoschema [34], which requires intrinsic temporal understanding capabilities. This highlights the gap in adapting web video-text pretrained models to more comprehensive video understanding tasks.
We attribute this gap to two primary factors: 1) The intrinsic complexity of the video modality. Videos introduce intrinsic complexities in both spatial and temporal dimensions, which are absent in static images. These complexities require extensive training on larger-scale data to effectively capture video dynamics. Furthermore, representing videos typically involves processing multiple frames, significantly increasing computational demands compared to image modeling. The dual challenges of large-scale training and increased computational requirements make video-language modeling particularly challenging. 2) The limitations of web language supervision. The language supervision in recent web video-text datasets primarily comes from subtitles associated with the videos. These subtitles are often frame-level descriptions that lack the necessary temporal associations crucial for video understanding. Moreover, this form of descriptive supervision is inefficient in building robust reasoning capabilities, especially in terms of temporal reasoning. This mismatch between the richness of video content and the limited language supervision hinders effective video-language modeling.
In this paper, we propose an innovative approach to develop video understanding capabilities using LLMs to simulate video dynamics. Instead of directly aligning video with language, we introduce a textual video representation — a sequence of textual frames designed to mimic real visual dynamics. This textual video can be readily generated by advanced LLMs and effectively simulates various video dynamics by describing them in text. Specifically, we present a Textual Video (TextVid) dataset, automatically generated by LLMs. TextVid includes: 1) Textual videos (hereinafter referred to as “Tideo”), which consist of a sequence of textual frames crafted to mimic the keyframes of real videos, and 2) Tideo annotations, including comprehensive Tideo-level dense descriptions and varied question-answer (QA) pairs. These annotations are of high quality and closely align with the Tideo content, by virtue of the powerful capability of LLM in language generation.
Building on the proposed TextVid dataset, we introduce the Text-Only Pre-Alignment (TOPA) framework, to effectively and efficiently pre-align LLMs with the video modality, reducing the need for costly video-text pre-training. We introduce three tasks for video-LLM pre-alignment: Tideo summarization, Tideo QA and multi-choice Tideo QA. To bridge the gap between textual Tideos and visual videos, we leverage the CLIP model for feature extraction. Specifically, we employ the CLIP text encoder to extract frame-level representations for Tideos, and the CLIP visual encoder for real videos. During the text-only pre-alignment phase, the LLM learns to process continuous CLIP text features. In the real video inference phase, it transitions to handling continuous CLIP image features. Due to the aligned CLIP image-text feature space, the LLM can adapt to real video inputs despite being trained on textual representations. Our main contributions include:
(1) We propose a novel Text-Only Pre-Alignment (TOPA) framework to extend Large Language Models (LLMs) for video understanding. TOPA aligns LLMs with the video modality efficiently and effectively without the need for training on real videos, reducing the costs for video-text pre-training.
(2) We introduce TextVid, a textual video dataset automatically generated by advanced LLMs. TextVid dataset comprises 721K diverse Tideos along with associated high-quality annotations, which include detailed Tideo descriptions and a variety of question-answer pairs.
(3) Extensive experiments demonstrate TOPA’s effectiveness across various video understanding tasks. Particularly, the TOPA-Llama2-13B model achieves 51.0% Top 1 accuracy in the challenging EgoSchema benchmark, outperforming previous video-text pretraining methods and approaching recent GPT-4-based video agents.
2 Related Work
Vision-Language Alignment. Radford et al. [43] align the vision and language in a common feature space via contrastive learning with large-scale web image-text data. [1, 23, 29, 79] align the visual model with advanced LLM via training on image-caption pairs and interleaved image-text data. [7, 19, 28, 73] explore modeling video sequences within LLM spaces, leveraging LLM for video-language understanding. In this paper, we focus on video-LLM alignment. Rather than using multimodal data for vision-language alignment, we introduce a novel text-only pre-alignment framework to extend LLMs for video understanding without pre-training on real video-text data.
LLMs for multimodal data augmentation. Recent research explores to leverage LLMs to enhance the multimodal data. A line of work [5, 12, 29] use LLMs for refining captions or extending the image caption pairs to diverse visual tasks like visual conversation and image editing. Another line of work [24, 32, 33, 42] further employ advanced LLM to enrich web video supervision for video instruction tuning. In this paper, rather than enhancing multimodal datasets, we propose generating text-only data consisting of "textual videos" and diverse language supervision, which aims to simulate real videos and their corresponding annotations.
Long-form Video Understanding. Long-form video understanding presents significant challenges due to the intricate spatial and temporal dynamics. Conventional video-text pretraining approaches [4, 40, 57, 58, 67, 80] utilize extensive web video-caption data for video-language alignment. Another line of research [18, 48, 71] seeks to adapt recent image MLLMs to video understanding. Recent research [24, 58, 73, 78] employ video instruction-tuning for video-LLM alignment to enhance video-language understanding. Recently, [9, 37, 49, 53, 55, 68, 72] combine the LLM with various VLM tools as video agents to perform video-understanding tasks. In this paper, we propose a novel text-only pre-alignment framework to efficiently and effectively align LLMs with videos without pre-training on real videos.
Figure 1: Overview of the proposed Text-Only Pre-Alignment (TOPA) framework. Left: The pipeline used for generating the TextVid dataset. Right: The video-LLM alignment framework. During text-only pre-alignment, the LLM learns to process continuous CLIP text features. In zero-shot inference, the LLM uses projected CLIP visual features as input. Additionally, TOPA supports supervised fine-tuning on downstream video datasets to further improve the performance.
3 Method
In this section, we detail the TOPA framework. We first introduce the data generation pipeline of TextVid Dataset (Section 3.1). Next, we describe how to align the Tideo representation with LLM (Section 3.2). Finally, we discuss adapting the text-only aligned video-LLM model for real video inference (Section 3.3). An overview is illustrated in Figure 1.
3.1 TextVid Dataset
This dataset, comprising textual videos (Tideos) and associated annotations, is generated by an advanced LLM (i.e., Gemini Pro 1.0 [50]). The data generation pipeline is detailed in Appendix B. Each Tideo is presented in a textual format and contains 5-15 sequential frames. Each frame includes a frame caption that describes the scene and multiple object captions. To enhance understanding and interaction with these Tideos, the dataset features a dense description summarizing the Tideo, as well as a set of multiple-choice questions and answers related to the Tideo content. The structure of each element is as follows:
Dataset Element: Tideo: Sequence of textual frames $\{T_{1},T_{2},\ldots,T_{n}\},5\leq n\leq 15$ For each frame $T_{i}$: Frame caption: $C_{i}$ Object captions: $D_{i,j}$ for main objects in $T_{i}$ Annotations: Global Dense Description of the Tideo: $D_{V}$ Set of Questions-Options-Answers: $\{(Q_{k},O_{k},A_{k})\}$
There are two major advantages of the TextVid dataset. (1) The large-scale and diverse Tideos. As the dataset is text-only and fully generated by an LLM, the size of TextVid is unlimited. Moreover, the Tideos can cover a broad range of domains by simply prompting the language model with appropriate conditions. It is distinctly different from previous web video-text dataset like Howto100M [35] that are limited to specific human-centric instructional videos. In practice, we enhance the diversity of TextVid by randomly sampling video captions from WebVid-2M [3], video titles from Howto100m [35], video tasks from Ego4D [14] and object names with descriptions from WordNet [36] as a condition of prompts. These varied prompts enable the language model to generate a diverse dataset. (2) The high-quality, consistent and free-form language supervision. The language supervisions are generated along with Tideos. The advanced capabilities of LLM ensure the quality of these supervisions, making them less noisy than web video-text data. Moreover, both the Tideo and the supervision are in textual format, making the supervision closely aligned with the Tideo’s content. Additionally, the format of the language supervision is unrestricted. For example, we prompt an LLM to generate dense description and multi-choice QA pairs as language supervision in TextVid.
3.2 Text-only Pre-alignment
Video-LLM alignment. The goal of video-LLM alignment is to extend the language-only LLMs for processing video inputs. Given a video sampled with $n$ frames $\{\mathbf{I}_{1},\mathbf{I}_{2},\ldots,\mathbf{I}_{n}\}$, we follow recent work [19, 67] to leverage a frozen CLIP model to extract the frame-level visual feature, formulated as $\mathbf{f}^{v}_{i}={E}_{\text{image}}(\mathbf{I}_{i})$, where $E_{\text{image}}$ denotes CLIP image encoder. The CLIP features are then projected into the LLM space via a simple linear layer, denoted as ${G}({P}(\mathbf{f}^{v}_{1}),...,{P}(\mathbf{f}^{v}_{n}))$, where ${G}$ denotes a language model and ${P}$ denotes a projection layer that projects the CLIP feature to LLM space.
Tideo Representation. In this work, we leverage Tideos (c.f. Section 3.1) for video-LLM pre-alignment instead of training on real videos. Specifically, given the textual frame $T_{i}$, we employ CLIP text encoder to extract the frame representation from frame caption $C_{i}$ and detailed object captions $D_{i}$, represented as $\mathbf{f}^{t}_{i}={F}_{\text{fusion}}({E}_{\text{text}}(C_{i}),{E}_{\text{text}}(D_{i,1}),...,{E}_{\text{text}}(D_{i,j}))$, where $F_{\text{fusion}}$ is a fusion function such as simple average pooling, and $E_{\text{text}}$ denotes the CLIP text encoder. A Tideo with $n$ textual frames is represented as $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{n}\}$
Text-only Training. Given the Tideo $T$, dense Tideo-level description $D_{V}$, and QA pairs with multiple choices $\{(Q_{k},O_{k},A_{k})\}$, we introduce the following tasks for Tideo-LLM alignment: (1) Tideo Summarization: Given the Tideo, generate a detailed description to summarize the Tideo; (2) Tideo QA: Given the Tideo and question, predict the answer; (3) Multi-choice Tideo QA: Given the Tideo, question and multiple choices, choose the correct answer from the candidates. We employ a unified auto-regressive Language Modeling (LM) objective for these three tasks:
$\mathcal{L}_{\text{LM}}(\theta_{{G}},\theta_{P})=-\frac{1}{|t|}\sum_{i=1}^{|t|}\log{G}\big{(}t_{i}|{P}(\mathbf{V}),{Z},t_{<i}\big{)},$ (1)
where $\mathbf{V}$ denotes the Tideo representation, and $\mathbf{V}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{i}\}$ during the text-only training, Z denotes the task specific condition tokens and $t_{i}$ denotes the $i_{th}$ target token. $\theta_{{G}}$ and $\theta_{{P}}$ denote the learnable parameters of the LLM adapter and the projection layer $P$, respectively. In practice, we use the following format as the LLM input: {Task Instruction}. Video:$\mathbf{f}^{t}_{1}$, …, $\mathbf{f}^{t}_{i}$}. {Task Conditions}. Answer: {Predict Targets}. For the Tideo summarization task, the target is detailed Tideo descriptions. For Tideo QA task, the target is the answer and the condition is the question. For multi-choice Tideo QA task, the target is the correct option and the condition consists with question and options. The details of the task-specific prompts are included in Appendix D.1.
3.3 Zero-shot Inference
Section 3.2 introduces the text-only pre-alignment using the TextVid dataset. In this section, we detail how to adapt this text-only pre-aligned LLM for real videos. During pre-alignment, we leverage the textual representation $\mathbf{V}^{t}=\{\mathbf{f}^{t}_{1},...,\mathbf{f}^{t}_{i}\}$ as the Tideo representation. During inference, we take real videos features as input, i.e., $\mathbf{V}^{v}=\{\mathbf{f}^{v}_{1},...,\mathbf{f}^{v}_{i}\}$, where $\mathbf{f}^{v}_{i}=E_{\text{image}}(\mathbf{I}_{i})$. These two modality features $\mathbf{f}^{t}$ and $\mathbf{f}^{v}$ that come from CLIP image encoder and CLIP text encoder are aligned via CLIP pre-training. This aligned image-text representation makes it possible to perform zero-shot inference without additional finetuning. However, the modality gap phenomenon [15, 26, 27, 39, 75], i.e., CLIP image feature and CLIP text feature are located in two completely separate regions of the feature space, prevents us from directly taking the visual feature $\mathbf{f}^{v}$ as the textual feature $\mathbf{f}^{t}$. To bridge this modality gap, we follow DeCap [26] to employ a support memory to project the CLIP visual feature into the CLIP text feature space. This training-free projection process is formulated as:
$\mathbf{f}^{v\rightarrow t}=\sum_{i=1}^{N}w_{i}*\mathbf{m}_{i}=\sum_{i=1}^{N}\frac{\exp((\mathbf{m}_{i}^{\top}\mathbf{f}^{v})/\tau)}{\sum_{k=1}^{N}\exp((\mathbf{m}_{k}^{\top}\mathbf{f}^{v})/\tau)}*\mathbf{m}_{i},$ (2)
where $\mathbf{m}_{i}$ denotes CLIP text feature from a pre-constructed memory of size $N$, $\mathbf{f}^{v}$ denotes input frame feature of real video and $\mathbf{f}^{v\rightarrow t}$ denotes the projected feature. During zero-shot inference, we take the $\mathbf{V}^{{v\rightarrow t}}=\{\mathbf{f}^{v\rightarrow t}_{1},...,\mathbf{f}^{{v\rightarrow t}}_{i}\}$ as the real video’s representation.
3.4 Implementation Details
We leverage Llama2-7B, Llama2-13B [51] and Llama3-8B as the LLM backbone. Additionally, we employ the Llama-adapter [74] with an adaptation embedding length of 50 for efficient finetuning. We utilize CLIP-ViT-L as the multimodal encoder. We employ a simple linear layer to project the CLIP feature into the LLM feature space. The CLIP model and LLM backbone are frozen. The projection layer and additional Llama-adapter are trainable. For text-only pre-alignment, we uniformly sample the Tideos into 10 frames. We train the model on a mixture of tasks comprising Tideo summarization, Tideo QA, multi-choice Tideo QA with the ratio of 1:1:2. TOPA-Llama2-7B and TOPA-Llama3-8B are trained on four 40G-A100 GPUs in one day. TOPA-Llama2-13B is trained in two days. For zero-shot inference, we construct a memory for cross-modal projection, consisting of 2M CLIP text features sampled from captions in the TextVid dataset. More training details of TOPA and baselines are included in Appendix C.2.
4 Experiments
TOPA enables the LLM to perform various video understanding tasks as shown in Figure 2. In this section, we evaluate TOPA on multi-choice video QA and video captioning tasks. Section 4.1 evaluates TOPA on NeXT-QA [62], STAR [61], TVQA [20], and recent challenging EgoSchema [34] benchmarks with the zero-shot setting. We further evaluate TOPA on multi-choice video QA with the finetuning setting (Section 4.2) and zero-shot video captioning task (Section 4.3). In Section 4.4, we conduct ablation study on the LLM prior and input video frames. We mainly compare TOPA against four categories of video understanding methods as follows:
(1) Web video pre-training approaches [4, 40, 56, 58, 67]. This line of work aims to develop general video-language models by leveraging extensive web videos, using associated video captions or audio as weak supervision signals.
(2) Adapt image MLLMs for video understanding [18, 71]. These approaches aim to extend the image understanding capabilities of recent vision-language models (VLMs) to video understanding. Specifically, SeViLa [71] utilizes BLIP-2 for localizing and understanding key frames of a video. IG-VLM [18] converts video into a composite image by arranging the video frames into a grid layout.
(3) Instructed Video MLLMs [24, 58, 73, 78]. Recent research employ video instruction-tuning for video-LLM alignment to enhance video-language understanding.
(4) LLM-based video agents [37, 52, 53, 55, 72]. This line of work leverages LLMs like GPT-3.5 and GPT-4 as an agent to understand a video by designing and executing a series of actions. The language-only agents perceive visual information via recent foundation VLMs (e.g., CLIP [43], BLIP-2 [23], LaViLa [77] and PALI [8]).
(5) Our text-only pre-alignment. Different from the above works, TOPA leverages the proposed TextVid dataset for video-LLM pre-alignment, enabling the LLM to process continuous features. Thus, performing video understanding tasks.
4.1 Zero-Shot Evaluation on Multi-Choice Video QA
In Section 4.1.1, we compare TOPA with a wide range of approaches on a recent challenging long-form video understanding benchmark, i.e., EgoSchema [34]. In Section 4.1.2, we further evaluate TOPA on conventional multi-choice video QA benchmarks including NeXT-QA [62], STAR [61] and TVQA [20]. In Section 4.1.3, we evaluate TOPA on recent comprehensive video understanding benchmark, i.e., MVBench [25]. The details of these benchmarks are included in Appendix C.1.
4.1.1 Zero-shot Results on EgoSchema
Table 1: Zero-shot results on Egoschema [34] full set. Methods that leverage closed-source LLMs are marked in gray. $\dagger$ denotes the model is trained with in-domain egocentric videos from Ego4D [14].
denotes results on EgoSchema subset. Results of InterVideo and FrozenBiLM are sourced from  [34]. Results of SeViLA are sourced from [40].
  Core VLMs Core LLMs Acc@1
Human Eval [34][NeurIPS 2023] - - 75.0%
Gemini-1.5-Pro [44][arXiv 2024.2 ] - Gemini-1.5-Pro 63.2%
(Pre-train on web video-text data)
FrozenBiLM [67][NeurIPS 2022] - - 26.9%
InterVideo [56][arXiv 2022.12] - - 32.1%
LongViViT [40][CVPR  2024 ] - - 33.3%
MC-ViT-L [4][ICML  2024 ] - - 44.4%
InterVideo2s3-6B [58][arXiv 2024.3 ] - - 41.1%
(Adapt image MLLMs for video understanding)
SeViLA [71][NeurIPS 2023] BLIP-2 FLAN-T5-XL [10] 22.7%
IG-VLM [18][arXiv 2024.3 ] LLaVA v1.6-7B Vicuna-7B 35.8%
IG-VLM [18][arXiv 2024.3 ] LLaVA v1.6-13B Vicuna-13B 47.0%
(LLM-based video agents)
Vamos [53][arXiv 2023.11] BLIP-2 Llama2-13B 36.7%
Vamos [53][arXiv 2023.11] BLIP-2 GPT-3.5 41.2%
Vamos [53][arXiv 2023.11] BLIP-2 GPT-4 48.3%
LLoVi [72][arXiv 2024.3 ] LaViLa GPT-3.5 50.3%
VideoAgent [55][arXiv 2024.3 ] LaViLa GPT-4 54.1%
MoReVQA [37][arXiv 2024.4 ] PALI-3-5B [8] PaLM-2 [2] 51.7%
(Our Text-Only Pre-Alignment) TOPA CLIP-ViT-L Llama2-7B 41.2%
TOPA CLIP-ViT-L Llama3-8B 44.2%
TOPA CLIP-ViT-L Llama2-13B 51.0%
Results on Egoschema. Table 1 shows the results. We compare our method against a range of recent approaches in video understanding. Our proposed text-only pre-alignment framework, despite training without real videos, shows impressive results on the Egoschema benchmark. TOPA outperforms previous image-based adaptation approach IG-VLM and video agents LLoVi and Vamos with the same scale LLM (Llama2-7B and Llama2-13B). Moreover, TOPA-Llama2 demonstrates consistent improvements when scaled up with a larger LLM backbone, indicating the effectiveness of LLMs in complex video-language understanding scenarios.
Discussion 1: The necessity of high-quality language supervision for video understanding. Recent video pre-training approaches like LongViVit [40] and InternVideo [56], despite training on million-level web video-text data, show inferior performance on Egoschema evaluation. These results highlight the inefficacy and inefficiency of conventional contrastive pre-training in understanding long-form videos, primarily due to noisy and simplistic language supervision. In contrast, our TOPA, trained on 721K virtual video (i.e., Tideo) with high-quality language supervision, shows impressive results on Egoschema. It indicates that, unlike image understanding which significantly benefits from leveraging web language as supervision, video understanding may require more precise and accurate language supervision to better capture the complex visual dynamics.
Discussion 2: Video agents versus end-to-end video-LLM modeling. Video agents have shown impressive results on the EgoSchema benchmark, aided by advanced LLMs and VLMs. However, a significant limitation of these approaches is their heavy reliance on the powerful LLMs. For example, the accuracy of Vamos drops by -11.6% when the GPT-4 is replaced with Llama2-13B. The reliance on powerful closed-source LLMs restricts its application fields and introduces external overheads. Additionally, the inference speed of these approaches is another concern, since it involves multiple interactions with both VLMs and LLMs. In contrast, end-to-end video-LLM models, which condense the video into a sequence of embeddings as the input of LLM, are more efficient. Moreover, video agents make decisions based on the language format clues collected by VLMs. Converting the video content into language clues may lead to a limited upper bound compared to end-to-end modeling.
Figure 2: Examples of TOPA-LLama2-13B for video-language understanding. Given a video, TOPA is able to summarize the video content and answer the questions.
4.1.2 Zero-shot Results on NExT-QA, STAR and TVQA
Table 2 shows the multi-choice video QA results across various benchmarks. TOPA achieves impressive performance on the TVQA and EgoSchema benchmarks, significantly outperforming previous video pre-training models and image-to-video adaptation approaches. This indicates that our TOPA framework effectively enables LLMs to handle video input, despite not being pre-trained on real videos. However, for the NeXT-QA and STAR benchmarks, TOPA underperforms compared to SeViLA and IG-VLM. A major reason is that these benchmarks involve many fine-grained visual questions, including those about object locations and relationships. SeViLA and IG-VLM, benefiting from the advanced image-understanding capabilities of pre-trained VLMs such as LLaVA, excel in answering these fine-grained visual questions. In contrast, our TOPA framework primarily focuses on high-level semantic alignment. Moreover, during zero-shot inference, we project the visual features into the text feature space to bridge the modality gap, as described in Eq. 2. This cross-modal semantic projection process tends to overlook fine-grained visual details, such as object locations, which leads to inferior performance on the STAR benchmark. We provide extensive qualitative results to illustrate TOPA’s advantages and limitations across various video understanding tasks in Appendix A.3.
Table 2: Zero-shot results on multi-choice video QA benchmarks.
 Model (# Frames) NExT-QA STAR   TVQA EgoSchema
Tem. Cau. Des. Avg. Int. Seq. Pre. Fea. Avg.
FrozenBiLM (10) [67] - - - - - - - - - 29.7 26.9
InternVideo (8) [56] 43.4 48.0 65.1 49.1 43.8 43.2 42.3 37.4 41.6 35.9 32.1
SeViLA (32 $\rightarrow$ 4) [71] 61.3 61.5 75.6 63.6 48.3 45.0 44.4 40.8 44.6 38.2 22.7
IG-VLM-Llava7B (6) [18] 63.1 57.3 74.9 63.1 49.3 50.1 48.4 48.8 49.6 42.1 35.8
IG-VLM-Llava13B (6) [18] 61.6 55.7 70.8 61.2 51.5 52.0 51.0 51.8 51.7 44.5 47.0
TOPA-Llama2-7B (10) 53.4 61.3 68.3 59.9 36.4 45.6 39.3 36.3 41.3 48.2 41.2
TOPA-Llama3-8B (10) 53.0 61.9 64.5 59.5 40.8 43.1 39.4 34.5 41.4 48.5 44.2
TOPA-Llama2-13B (10) 57.2 63.6 68.9 62.1 41.6 46.2 44.2 36.7 43.0 50.2 51.0
4.1.3 Zero-shot Results on MVBench
MVBench [25] is a recent video-language understanding benchmark that covers 20 challenging video tasks, regrouped from existing video-language benchmarks. For details of these 20 video understanding tasks, please refer to  [25]. Table 3 shows the results. TOPA demonstrates impressive results compared to previous image MLLMs and video MLLMs. It excels particularly in tasks requiring high-level video-language understanding, such as Scene Transition (ST), Episodic Reasoning (ER), and Unexpected Action (UA). Surprisingly, TOPA also excels in the Action Localization (AL) task, which requires identifying the moment an action occurs. This indicates that the text-only pre-alignment enables the LLM to understand visual temporal sequences. However, TOPA struggles with tasks that demand fine-grained visual understanding, such as Moving Direction (MR), Action Antonym (AA), and Object Shuffle (OS). A common challenge in these tasks is the requirement for detailed visual understanding. For example, Action Antonym involves identifying the direction of an action, while Object Shuffle involves locating objects. TOPA struggles in these fine-grained visual tasks since it is trained with CLIP text features. The modality gap between CLIP text features and image features hinders TOPA from capturing visual details. Further video instruction tuning might address this limitation, which we leave for future work. We provide qualitative results in Figure 4 - 7 to illustrate TOPA’s advantages and limitations on various video understanding tasks.
Model LLM Avg AS AP AA FA UA OE OI OS MD AL ST AC MC MA SC FP CO EN ER CI
Random - 27.3 25.0 25.0 33.3 25.0 25.0 33.3 25.0 33.3 25.0 25.0 25.0 33.3 25.0 33.3 33.3 25.0 33.3 25.0 20.0 30.9
Image MLLMs: Following  [11], all models take 4 frames as input, with the output embeddings concatenated before feeding into the LLM.
mPLUG-Owl-I [64] LLaMA-7B 29.4 25.0 20.0 44.5 27.0 23.5 36.0 24.0 34.0 23.0 24.0 34.5 34.5 22.0 31.5 40.0 24.0 37.0 25.5 21.0 37.0
BLIP2 [23] FlanT5-XL 31.4 24.5 29.0 33.5 17.0 42.0 51.5 26.0 31.0 25.5 26.0 32.5 25.5 30.0 40.0 42.0 27.0 30.0 26.0 37.0 31.0
LLaMA-Adapter [74] LLaMA-7B 31.7 23.0 28.0 51.0 30.0 33.0 53.5 32.5 33.5 25.5 21.5 30.5 29.0 22.5 41.5 39.5 25.0 31.5 22.5 28.0 32.0
Otter-I [21] MPT-7B 33.5 34.5 32.0 39.5 30.5 38.5 48.5 44.0 29.5 19.0 25.5 55.0 20.0 32.5 28.5 39.0 28.0 27.0 32.0 29.0 36.5
MiniGPT-4 [79] Vicuna-7B 18.8 16.0 18.0 26.0 21.5 16.0 29.5 25.5 13.0 11.5 12.0 9.5 32.5 15.5 8.0 34.0 26.0 29.5 19.0 9.9 3.0
InstructBLIP [11] Vicuna-7B 32.5 20.0 16.5 46.0 24.5 46.0 51.0 26.0 37.5 22.0 23.0 46.5 42.5 26.5 40.5 32.0 25.5 30.0 25.5 30.5 38.0
LLaVA [29] Vicuna-7B 36.0 28.0 39.5 63.0 30.5 39.0 53.0 41.0 41.5 23.0 20.5 45.0 34.0 20.5 38.5 47.0 25.0 36.0 27.0 26.5 42.0
Video MLLMs: All models take 16 frames as input.
Otter-V [21] LLaMA-7B 26.8 23.0 23.0 27.5 27.0 29.5 53.0 28.0 33.0 24.5 23.5 27.5 26.0 28.5 18.0 38.5 22.0 22.0 23.5 19.0 19.5
mPLUG-Owl-V [69] LLaMA-7B 29.7 22.0 28.0 34.0 29.0 29.0 40.5 27.0 31.5 27.0 23.0 29.0 31.5 27.0 40.0 44.0 24.0 31.0 26.0 20.5 29.5
Instructed Video MLLMs: All models take 16 frames as input, with the exception of VideoChatGPT, which uses 100 frames.
VideoChatGPT [33] Vicuna-7B 32.7 23.5 26.0 62.0 22.5 26.5 54.0 28.0 40.0 23.0 20.0 31.0 30.5 25.5 39.5 48.5 29.0 33.0 29.5 26.0 35.5
VideoLLaMA [73] Vicuna-7B 34.1 27.5 25.5 51.0 29.0 39.0 48.0 40.5 38.0 22.5 22.5 43.0 34.0 22.5 32.5 45.5 32.5 40.0 30.0 21.0 37.0
VideoChat [24] Vicuna-7B 35.5 33.5 26.5 56.0 33.5 40.5 53.0 40.5 30.0 25.5 27.0 48.5 35.0 20.5 42.5 46.0 26.5 41.0 23.5 23.5 36.0
VideoChat2 [25] Vicuna-7B 51.1 66.0 47.5 83.5 49.5 60.0 58.0 71.5 42.5 23.0 23.0 88.5 39.0 42.0 58.5 44.0 49.0 36.5 35.0 40.5 65.5
Text-Only Pre-Alignment Video-MLLM: TOPA zero-shot inference with 10 frames.
TOPA-ZeroShot LLama2-7B 39.8 42.0 38.5 35.0 34.5 66.0 52.5 47.5 28.0 22.0 37.5 81.0 38.0 24.0 42.5 41.5 28.5 34.0 23.5 49.0 30.5
TOPA-ZeroShot LLama2-13B 42.5 38.0 40.0 42.5 35.0 69.0 52.5 58.5 29.5 22.5 43.5 80.5 38.0 25.5 43.0 43.0 29.5 37.5 38.5 50.0 32.5
Table 3: Evaluation results on MVBench. The results of other approaches are sourced from  [25]. We gray out the results of VideoChat2 because it utilizes extensive manually annotated video data from downstream video benchmarks for video instruction tuning.
AS AP AA FA UA OE OI OS MD AL
Action Action Action Fine-grained Unexpected Object Object Object Moving Action
Sequence Prediction Antonym Action Action Existence Interaction Shuffle Direction Localization
STAR [61] STAR [61] PAXION [59] MiT V1 [38] FunQA [63] CLEVRER [70] STAR [61] Perception Test [41] CLEVRER [70] Charades-STA [13]
ST AC MC MA SC FP CO EN ER CI
Scene Action Moving Moving State Fine-grained Character Egocentric Episodic Counterfactual
Transition Counting Counting Attribute Change Pose Order Navigation Reasoning Inference
MovieNet [16] Perception Test [41] CLEVRER [70] CLEVRER [70] Perception Test [41] NTU RGB+D [30] Perception Test [41] VLN-CE [60] TVQA [20] CLEVRER [70]
4.2 Finetuning Results
In this section, we further finetune the pre-aligned TOPA models to study the benefits of TOPA for downstream supervised learning. During finetuning, TOPA directly takes the video feature as input without the cross-modal projection. More finetuning details for each dataset are provided in Appendix C.2. Table 4 shows the finetuning results on multi-choice video QA dataset including NExT-QA, STAR and TVQA. For comparison, we include baseline models without text-only pretraining. Our text-only pre-alignment consistently improves the performance across three benchmarks. Notably, TOPA-Llama2-7B achieves 67.1% accuracy on TVQA, outperforming other approaches by a large margin. Additionally, in the zero-shot evaluation (Table 2), TOPA-Llama2-7B shows inferior results on STAR compared to SeViLA, due to its limitation in handling fine-grained visual details. However, in this finetuning setting, TOPA-Llama2-7B outperforms SeViLA, indicating that finetuning on downstream video data effectively mitigates TOPA’s limitations in processing fine-grained visual details. These results suggest that our text-only pre-alignment, despite not training with real video data, has a similar effect to conventional video-text pre-training.
Table 4: Finetuning results on NExT-QA, STAR and TVQA.
 Model (# Frames) NExT-QA STAR TVQA
Tem. Cau. Des. Avg. Int. Seq. Pre. Fea. Avg.
FrozenBiLM (10) [67] - - - - - - - - - 57.5
InternVideo (8) [56] 58.5 62.5 75.8 63.2 62.7 65.6 54.9 51.9 58.7 57.2
[…]
SeViLA (32 $\rightarrow$ 4) [71] 69.4 74.2 81.3 73.8 63.7 70.4 63.1 62.4 64.9 61.6
Llama-VQA-7B (10) [19] 69.2 72.7 75.8 72.0 66.2 67.9 57.2 52.7 65.4 -
Baseline (10) 65.3 69.0 72.6 68.4 60.8 61.5 49.2 49.8 59.4 63.8
TOPA-Llama2-7B (10) 71.3 74.2 78.5 73.9 66.8 68.9 59.1 55.5 66.4 67.1
+6.0 +5.2 +5.9 +5.5 +6.0 +7.4 +9.9 +5.7 +7.0 +3.3
[…]
TOPA-Llama3-8B (10) 70.1 74.5 74.6 73.1 66.3 67.0 59.1 56.5 65.4 68.1
+4.1 +4.8 +0.9 +4.0 +4.9 +4.6 +8.5 +4.7 +5.1 +1.9
[…]
TOPA-Llama2-13B (10) 72.1 75.8 79.3 75.1 66.8 68.3 61.0 55.1 66.3 69.0
+4.3 +4.2 +4.1 +4.2 +8.1 +8.8 +6.7 +3.3 +8.1 +2.4
Data Efficient Finetuning Figure 3 shows the results of finetuning LLMs with various ratios of training data. TOPA trained with 10% data achieves 64.7% Top 1 accuracy on NeXT-QA benchmark, significantly outperforming the baseline that without text-only pre-alignment. Besides, when trained with less than 20% data, the baseline model even performs worse than TOPA-ZS on NeXT-QA and TVQA, demonstrating the effectiveness of TOPA in limited annotated data scenarios.
Figure 3: Results of finetuning TOPA with various ratios of training data.
4.3 Video Captioning
Results on Zero-shot Video Captioning. We further perform zero-shot video captioning on MSR-VTT [65] and VATEX [54]. As shown in Table 5, TOPA largely outperforms previous text-only approaches like Decap which is trained on captions sourced from CC3M [47]. TOPA even outperforms the video-text pre-training approaches like VideoCoCa, which is pre-trained on millions of videos-text data, demonstrating that TOPA is an efficient and effective framework for video-LLM alignment.
[…]
  Training data MSR-VTT VATEX
(Web video-text Pre-training)
VideoCoCa-g [66] 144M VT 27.1 22.8
Flamingo-3B  [1] 27M VT & 2.1B IT & 43M WP - 40.1
Flamingo-9B  [1] 27M VT & 2.1B IT & 43M WP - 39.5
VideoPrism-B [76] w/ PaLM-2-1B 618M VT 40.3 24.2
VideoPrism-B [76] w/ PaLM-2-8B 618M VT 38.5 31.7
(Text-only Pre-training)
Decap [26] 3M Captions 18.6 18.7
TOPA-Llama2-7B 721K TextVid 32.9 31.0
TOPA-Llama2-13B 721K TextVid 33.4 32.0
4.4 Ablations
LLM Prior in Video-language Understanding. To investigate the impact of LLM prior in multi-choice video QA, we conduct experiments on EgoSchema with the blind setting, where only the questions and choices are provided to the LLM. Table 7 shows the results. Bard and GPT-4-Turbo achieve 33.2% and 30.8% accuracy, respectively. Gemini-Pro-1.0 reaches 38.2% accuracy. These blind results of advanced LLMs suggest that in some video QA cases, LLMs can accurately choose the correct answer solely based on the question and choices, without visual input. However, the blind performance of Llama2-7B and Llama2-13B is inferior, potentially due to their smaller model size. After training on the TextVid dataset, TOPA-Llama2-13B achieves a blind accuracy of 37.5% (or +11.7%), closely approaching that of Gemini-Pro-1.0 model. These results suggest that text-only pre-alignment can effectively prepare LLMs for downstream video-language tasks by leveraging specialized text-only tasks, even in complex scenarios where the original LLMs are limited.
[…]
Table 6: Blind results on EgoSchema. $\dagger$ denotes results sourced from [4].
Table 7: Ablation on video frames.
The Impact of Video Frames. To better investigate TOPA’s capability in understanding temporal dynamics of real videos, we conduct experiments with different number of frames. Table 7 shows the results. Multiple frames input consistently enhances performance on NeXT-QA and Egoschema for both TOPA-Llama2-7B and TOPA-Llama2-13B. This indicates that the text-only pre-alignment effectively enables the LLM to handle multiple video frames, despite not being trained on real videos.
5 Limitations.
Modality gap in CLIP. Despite the fact that TOPA achieves impressive results. A significant limitation in TOPA is the gap between the CLIP text feature and CLIP image feature. On the one hand, we use the CLIP text feature for pre-alignment, while inference is with the CLIP visual feature. The modality gap makes the performance degrades, despite we employ a modality projection mechanism to mitigate it. On the other hand, the CLIP text features cannot fully capture the fine-grained visual details present in actual images, such as object locations and relationships. This limitation causes TOPA to struggle in scenarios where questions involve detailed visual information, as shown in Appendix A.3.
Struggles in fine-grained visual understanding. In TOPA, we propose textual videos to mimic real videos. However, this approach primarily focuses on keyframes understanding, which is insufficient for scenarios requiring the model to process hundreds of frames at high fps, such as action counting tasks. Besides, for the fine-grained action understanding scenarios, TOPA is unable to capture the fine-grained visual information. For example, in a scene where a person climbs a ladder, it is difficult for TOPA to identify whether the person is going up or down due to the limited capability to capture detailed visual dynamics. Further enhancing TOPA with video instruction tuning might address these limitations which we leave for future work.
6 Broader Impact
Academic Impact. TOPA’s methodology, which frees the need for costly video-text data collection and large-scale pre-training, lowers the barriers to entry for research and development in video-language understanding technologies. The text-only learning framework of TOPA may inspire researchers with limited resources to engage in cutting-edge multi-modal research, providing a more diverse range of perspectives and contributions to this field.
Social Impact. The ultimate objective of TOPA is to develop a general video-language understanding model. Its primary application enables users to extract information from long-form videos without the need for detailed viewing. Moreover, these capabilities for interpreting and managing video content could significantly enhance content moderation systems. Platforms hosting user-generated content could employ sophisticated video-language models to efficiently detect and mitigate the effects of inappropriate or harmful video content.
7 Conclusions
In this paper, we introduce TOPA, a text-only pre-alignment framework designed for aligning LLMs with video modality without requiring training on real videos. TOPA has demonstrated remarkable performance on the recent, challenging long-form video understanding benchmark, i.e., EgoSchema, showcasing that a text-only approach is effective in capturing the dynamics of long-form videos. Our approach, which includes data generation and text-only pre-alignment, has potential applications across various vision-language tasks where obtaining paired vision-language data is difficult.
Appendix
In Appendix A, we provide additional experiments and analysis.
  • In Appendix A.1, we further discuss the Multi-choice QA task and study the impact of the multi-choice Tideo QA pre-training.
  • In Appendix A.2, we study the impact of cross-modal projection (Eq. 2).
  • In Appendix A.3, we provide extensive qualitative results to illustrate TOPA’s advantages and limitations across various video understanding tasks.
Appendix B: The details of proposed TextVid dataset.
Appendix C.1, The details of benchmarks.
Appendix C.2, The training details of TOPA.
Appendix D: The prompts used in this paper.
Appendix E: Examples from TextVid.
Appendix A Additional Experiments
A.1 Further Discussion on Multi-Choice Video QA Task
Table 8: Multi-choice video QA on EgoSchema subset and full set. “Gap” refers to the difference in performance between the subset and the full set
 Method Eval Mode ES Subset ES Full Gap
LongViViT [40] Similarity 56.8 33.3 -23.5
MC-ViT-L [4] Similarity 62.6 44.0 -18.6
VideoAgent (GPT-4) [55] LLM Selection 60.2 54.1 -6.1
TOPA-LLama2-13B LLM Selection 51.2 51.0 -0.2
TOPA-LLama2-13B LLM Logits 67.5 41.6 -25.9
TOPA-LLama2-7B LLM Selection 40.4 41.2 +0.8
TOPA-LLama2-7B LLM Logits 64.5 41.7 -22.8
TOPA-LLama2-7B (w/o multi-choice training) LLM Selection 24.3 24.7 +0.4
TOPA-LLama2-7B (w/o multi-choice training) LLM Logits 65.1 40.5 -24.6
A significant advantage of the text-only framework is that we can utilize the LLM to automatically generate diverse language-based supervisions as needed, such as the multi-choice QA pairs. To explore the impact of the multi-choice QA training tasks, we conduct an ablation study as shown in Table 8. We would like to first introduce the different evaluation modes for multi-choice video QA tasks: (1) LLM Selection: Asking the LLM to predict the correct answer given the video-question-choices. (2) LLM Logits: Given the video and question as LLM context, we calculate the logits for each choice by averaging the logits of all words within the choice. The choice with higher logit tends to match the video-question context better and is thus selected as the predicted answer. (3) Similarity Comparison [40, 4]: Mapping the multiple question-choice pairs and video to a common feature space and calculating the similarity between the video and each question-choice.
The performance gap between the EgoSchema subset and full set. Previous work [40, 4] highlights a huge performance gap between the subset and the full set of EgoSchema as shown in Table 8. We observe that such a performance gap phenomenon also occurs in approaches based on LLM logits. However, but it diminishes or even disappears in methods employing LLM selection. We find that this phenomenon may be attributed to differences in the linguistic structures of the choices, as shown below. The choices in the subset often differ in several key works like “create”, “repair” and “clean”. The similarity or logit can effectively identify this keyword-level difference to select a more appropriate choice. Conversely, the choices in the full set display more substantial linguistic differences. These variations introduce significant language biases, i.e., some sentences naturally receive higher logits in LLM, complicating the reliance on similarity or logit for choice selection. In contrast, LLM selection methods take all the choices within the context, allowing the LLM to leverage its robust contextual understanding to select the correct choice.
Question-Choices examples from subset: Q: Can you summarize the primary objective and the steps the person took throughout the video to achieve it? ensure your answer captures the essence of the video without listing all actions. A: The main aim of the person’s primary objective was to create and build a new, sturdy wooden bench. B: The primary objective for the person was to thoroughly repair and restore the wooden bench. C: The person’s primary objective was to thoroughly clean and sanitize the wooden bench’s surface.
Question-Choices examples from full set: Q: Considering the entire video, what would you identify as the most crucial moments in the person’s shopping experience and why? A: Following a strict shopping list as a guideline and rejecting unfit produce. B: Conducting taste tests and checking for the freshness of each vegetable. C: Using math algorithm for optimal vegetable selection.
LLM for multi-choice QA. In Table 8, we observe a notable phenomenon where the TOPA models achieve impressive results on the subset with the logits evaluation mode. TOPA-LLama2-13B achieves 67.5% top1 accuracy, surpassing GPT-4-based video agents. However, when evaluated with the multi-choice selection mode, the performance of the subset declines to 51.2%, but the performance of the full set increases from 41.6% to 51.0%. These results suggest that while the LLM is capable of selecting the answer from multiple choices, it is less sensitive to the keywords within those choices. In contrast, the logit-based approach is sensitive to the keywords but has difficulty with complex sentence understanding.
The impact of the Multi-Choice Tideo QA pre-training. In Table 8, we report the results of TOPA without the multi-choice Tideo QA task, i.e., trained with Tideo summarization and Tideo QA tasks. In this case, we find that TOPA-LLama2-7B maintains similar performance when evaluated with the logit mode. However, there is a significant performance drop when evaluated with the multi-choice selection mode. This result suggests that while the LLM is adapted to process video inputs, its capability is somewhat constrained and can not extend to more complex video-language tasks beyond the pre-training tasks. This finding highlights the advantage of our text-only data generation and text-only pre-alignment framework, which enable us to develop a variety of pre-alignment tasks to better equip the LLM for general video-language tasks such as dense captioning, multi-choice video QA, and video chat.
A.2 The CLIP Modality Gap
TOPA is pretrained with CLIP text features while inferenced with CLIP image features. We employ a modality projection approach, i.e., Eq. 2, to bridge this CLIP modality gap during zero-shot inference. Table 9 shows the impact of Eq. 2. TOPA shows inferior results when directly taking the visual feature as input due to the modality gap problem. The projection approach effectively alleviates such a modality gap problem without additional training.
Table 9: Ablation on the modality projection (Eq. 2). Results on EgoSchema full set.
 Model without Eq. 2 with Eq. 2
TOPA-LLama2-7B 30.6 41.2
TOPA-LLama2-13B 38.3 51.0
A.3 Qualitative Results and Analysis
We present qualitative results to illustrate the capabilities and limitations of TOPA across various video understanding tasks. Figure 8 shows qualitative results on the NExT-QA validation set. Figure 9 shows qualitative results on the EgoSchema subset. Figure 4 - 7 shows qualitative results on 20 video understanding tasks from MVBench.
Figure 4: Qualitative results on MVBench (Task 1-5). The tasks where TOPA performs well, average, or poorly are marked in green, blue, and red colors respectively.
Figure 5: Qualitative results on MVBench (Task 6-10).
Figure 6: Qualitative results on MVBench (Task 11-15).
Figure 7: Qualitative results on MVBench (Task 15-20).
Figure 8: Qualitative results on NeXT-QA. TOPA effectively performs complex video understanding tasks. Additionally, a failure case is also shown in the figure, i.e., in the last sample, TOPA failed to accurately count the number of people.
Figure 9: EgoSchema presents unique challenges compared to previous video benchmarks. The questions in EgoSchema are complex and demand advanced video capabilities, encompassing both recognition and reasoning skills.
Appendix B The details of proposed TextVid dataset
We utilize Gemini Pro 1.0 API for our data generation process. We prompt the LLM to create textual videos along with associated annotations. To ensure a diverse dataset that covers a wide of domains, we add condition prompts including different themes, video captions, video events, and the names of main objects. Specifically, we leverage video titles from Howto100M [35], video captions from WebVid2M [3], video events from Ego4D [14], and object from Wordnet [36] as conditions to generate diverse textual videos. For Ego4D condition, we ask the LLM to mimic an ego-centric video to further improve the diversity of the dataset. For each data generation, we prompt the LLM with the task prompt and one of the condition prompts as shown in Figure 12. The statistics TextVid are shown in Table 10. Additionally, we provide Wordcloud of TextVid in Figure 10. The examples of TextVid are shown in Appendix E.
Figure 10: Wordcloud of TextVid. The frame caption (left) and the dense video caption (right).
Table 10: Statistics of TextVid.
  TextVid
Generated by Gemini Pro 1.0
# textual videos 721K
# each condition Video Title-213k; Video Caption-183K; Video Scenarios-205K; Object-120K
# all QA pairs 3.5M
# each question types What-2.5M; Why-410K; How-287K;Others-254K
# all frames 4.4M
Avg. frame 6.13
Appendix C Experimental Setup
C.1 Benchmarks
EgoSchema [34] is a challenging long-form video understanding benchmark with 5000 multi-choice questions. The videos in EgoSchema are sourced from Ego-4D [14], with an average length of 3 minutes, distinct from previous benchmarks that focused on shorter, seconds-long videos. The questions in EgoSchema are manually curated to demand long temporal reasoning. We report results on EgoSchema full set.
NExT-QA [62] is a multi-choice video QA benchmark for causal and temporal reasoning, including 5,440 natural videos. The average length of video is 44 seconds. We report results on NExT-QA validation set, which contains 570 videos and 5,000 multiple-choice questions.
STAR [61] is a benchmark for situated reasoning. It contains 22K video clips with an average length of 44 seconds. There are 4 different question types in STAR: Interaction (Int.), Sequence (Seq.), Prediction (Pre.), and Feasibility (Fea.). We report results on STAR validation set.
TVQA [20] is a benchmark containing 21k video clips with an average length of 76 seconds. We report results on TVQA validation set without subtitles.
MVbench [25] is a reorganized benchmark containing 20 video understanding tasks. These tasks are sourced from STAR [61], PAXION [59], MiT Vi [38], FunQA [63], Perception Test [41], Charades-STA [13], MovieNet [16], NTU RGB+D [30], VLN-CE [60] and TVQA [20].
C.2 The details of training and evaluation.
[…]
Table 11: Training hyper-parameters.
  Model Training Dataset Epoch Effective Batchsize blr Optimizer
(bs, #GPUs, grad accu)
Pre-training TOPA-LLama2-7B TextVid 721K 20 18x4x4 5e-3 AdamW
TOPA-LLama2-13B 4x4x8 8e-3 weight decay 0.1
TOPA-LLama3-8B 14x4x8 5e-3 warm up 1 epoch
Fine-tuning TOPA-LLama2-7B NextQA 5 20x4x4 5e-3
STAR 5 1e-2
TVQA 5e-3
TOPA-LLama2-13B NextQA 6x4x16 2e-3 AdamW [31]
STAR 2e-3 weight decay 0.1
TVQA 5e-3 warm up 1 epochs
TOPA-LLama3-8B NextQA 20x4x4 1e-2
STAR 5 1e-2
TVQA 5e-3
Baseline LLama2-7B NextQA 10 20x4x4 1e-2
STAR 10 2e-2
TVQA 2e-2
LLama2-13B NextQA 6x4x16 1e-2 AdamW
STAR 10 2e-2 weight decay 0.1
TVQA 2e-2 warm up 2 epochs
LLama3-8B NextQA 20x4x4 2e-2
STAR 10 2e-2
TVQA 1e-2
Appendix D Prompts
D.1 Text-only Training Prompts.
We use the following prompts for Text-only training. The prompts are partially based on [19].
Tideo Multi-choice QA: Instruction: Choose the correct answer based on the video and question.
Video: $\mathbf{f}^{t}_{1}$,…,$\mathbf{f}^{t}_{i}$}.
 Question: {Question}.
Choices:
(A): {Option A}. (B): {Option B}. (C): {Option C}. (D): {Option D}. (E): {Option E}.
Answer: The correct choice is
{Correct Choice}.
Tideo QA: Instruction: Predict the answer based on the video and question.
Video: $\mathbf{f}^{t}_{1}$,…,$\mathbf{f}^{t}_{i}$}.
Question: {Question}.
Answer:
{Answer}.
Tideo Description: Instruction: Generate a dense description for the video.
Video: $\mathbf{f}^{t}_{1}$,…,$\mathbf{f}^{t}_{i}$}.
Description:
{Tideo Description}.
D.2 Prompt for Gemini Blind Evaluation
In Table 7, we use the following prompt for the blind evaluation of Gemini-Pro-1.0 on Egoschema. The prompt is based on [4].
Figure 11: The multi-choice QA prompts used for the blind evaluation of Gemini-1.0-Pro.
D.3 Prompts for Data Generation
The prompts for TextVid generation are shown in Figure 12.
Figure 12: The prompts used in TextVid generation.
Appendix E Examples of TextVid